# Jailbreaking Attacks on GPT-based Text-to-Image Models

### v1

@@ 3. $PC^{2}$ Design
3.1. Threat Model
In this study, we consider a realistic black-box threat model. The primary goal of an adversary is to circumvent the safety filters of a T2I model to generate politically sensitive contents (PSCs), which can be weaponized for disinformation or propaganda. We assume the adversary operates under black-box constraints, meaning they have no access to the T2I model’s internal parameters, gradients, training datasets, or the specific logic of the safety filters. The adversary interacts with the target model solely through standard interfaces, such as web-based clients (e.g., ChatGPT’s gpt-image-1 interface), rather than direct API access, as the gpt-image-1 API requires organization verification (identity validation) making publicly available web clients the most realistic interaction channel from an adversarial perspective. This threat model is highly practical as it reflects the most common real-world vector for T2I misuse.
3.2. Politically Sensitive Term Detection and IPDM
Figure 5. Identity-Preserving Descriptive Mapping (IPDM) description example.
@@ 3.2. Politically Sensitive Term Detection and IPDM
Politically sensitive term detection is performed through a multi-stage pipeline with two independent processing paths. First, named entity recognition (NER) is applied to the input prompt to identify political figures (e.g., Donald Trump in Figure 5) and objects (e.g., Nazi flag in Figure 5). For each identified political figure, a dedicated inference step based on a language model is used to determine the most relevant country. Separately, $PC^{2}$ extracts noun phrases that may be associated with a specific country or with politically or socially sensitive topics. The extracted noun phrases are then analyzed using a language model (GPT-4o in our implementation) to determine their association with a specific country and their political or social sensitivity. This design enables the identification of politically sensitive terms beyond surface-level keyword matching. The inferred country information from both processing paths is retained for subsequent metric computation and is not used to modify the prompt at this stage.
For each identified controversial term, the IPDM is applied to generate a corresponding description as shown in Figure 5. Replacing explicit terms with IPDM-based descriptions serves two purposes: (i) it bypasses keyword-based filters that rely on matching explicit terms, and (ii) it enables the subsequent translation phase to operate even for terms that lack direct translation equivalents in the target language. IPDM is implemented using a language model, where $PC^{2}$ instructs the model to generate a 1–2 sentence quiz-style description that implicitly conveys the concept without explicitly naming the original term (see Appendix B)
3.3. Translation
@@
There are two types of targets require the selection of a final language to construct the adversarial prompt: the IPDM descriptions and the base prompt. The IPDM descriptions are generated in the previous step (Section 3.2), while the base prompt corresponds to the original prompt in which controversial terms are replaced with placeholders (i.e., <NCC_PLACE_HOLDER_x>). These placeholders are later substituted with alphabetical indices (e.g., A, B, etc.), each mapped to its corresponding IPDM description. Both targets are initially translated into 72 languages (Appendix A).
@@
For translation, we employ GPT-4o to maximize translation accuracy. Nevertheless, even high-performing language models may produce hallucinations or translation errors, particularly for low-resource languages. To mitigate this issue, we apply a back-translation procedure: each translated output is translated back into English and compared with the original content using cosine similarity. $PC^{2}$ filters out translations whose back-translated similarity score falls below an acceptable threshold (cosine similarity <0.8<0.8), thereby removing unreliable translations from further consideration.
3.4. Metrics
@@
To select an appropriate language for each target, we devise four metrics. These metrics assess the potential level of political sensitivity in the translated output for each language and assign a corresponding score. Notably, the metrics largely depend on the originating term of the description and are therefore primarily applicable to IPDM descriptions rather than the base prompt. This design choice is justified because the base prompt is intentionally neutral, containing only placeholders combined with action verbs (e.g., <NCC_PLACE_HOLDER_1> holding <NCC_PLACE_HOLDER_2>). Accordingly, for the base prompt, we select the language by computing the cumulative sum of the language scores obtained from the associated IPDM descriptions.
@@ 3.4.1. Keyword Common Knowledge-based Metric
The keyword common knowledge–based metric measures how controversially a political input may be perceived, using publicly available historical associations between the input’s keywords and countries, then aggregating those signals into geopolitical sensitivity score. For each original prompt $x_{k}\in\mathcal{X}$ (with $k\in\{1,\dots,K\}$), as shown in Section 3.2, we first identify a set of political phrases $\mathcal{E}_{k}=\mathrm{Extract}(x_{k})$. Here, $\mathrm{Extract}(\cdot)$ returns political keywords (e.g., political figures, political objects) that can serve as Wikipedia queries.
For each extracted political keyword $e\in\mathcal{E}_{k}$, we retrieve relevant Wikipedia paragraphs using $\mathrm{WikiSearch}(\cdot)$ and merge all retrieved content into a single paragraph set $\mathcal{P}_{k}=\bigcup_{e\in\mathcal{E}_{k}}\mathrm{WikiSearch}(e)$. The retrieved Wikipedia text is stored as paragraphs, which serve as the atomic retrieval units in a knowledge database. This database can be dynamically expanded as new phrases are detected in subsequent inputs. Each paragraph $p\in\mathcal{P}_{k}$ is encoded into a dense vector representation $\mathbf{v}_{k,p}=f_{embed}(p)\in\mathbb{R}^{d}$ using a text-embedding model $f_{embed}(\cdot)$, where $d$ is the embedding dimension.
To measure political sensitivity with respect to geopolitical actors, we construct a country-specific prompt by concatenating a fixed prefix with the country name. For example, country-specific prompt $q_{i}=\mathrm{Conf}\|\mathrm{name}_{i}$, where $\mathrm{Conf}$ is the string “Conflict with” and $\mathrm{name}_{i}$ denotes the name of country $i\in\{1,\dots,N\}$. We embed each country-specific prompt as $\mathbf{u}_{i}=f_{embed}(q_{i})\in\mathbb{R}^{d}$. For a given input $x_{k}$, we compute a country-level geopolitical sensitivity score by taking the maximum cosine similarity between the country prompt embedding and any evidence paragraph embedding retrieved for that input:
@@
(3) $\mathbf{S}_{kc}(k,\ell)=\left(S^{kc}_{k,\ell}\right)_{\ell\in\mathcal{L}}\in\mathbb{R}^{|\mathcal{L}|},$
@@ 3.4.2. Country Common Knowledge-based Metric
We first enumerate a set of countries $\mathcal{C}=\{1,\dots,N\}$ and retrieve relevant Wikipedia paragraphs for each country name $\mathrm{name}_{i}$ using $\mathrm{WikiSearch}(\cdot)$. We merge the retrieved content into a country-specific paragraph set $\mathcal{P}_{i}=\mathrm{WikiSearch}(\mathrm{name}_{i})$.
The collected Wikipedia text is segmented into paragraphs, which serve as the atomic retrieval units in a knowledge database. This database can be dynamically expanded as new countries are added. Each paragraph $p\in\mathcal{P}_{i}$ is encoded into a dense vector representation $\mathbf{v}_{i,p}=f_{embed}(p)\in\mathbb{R}^{d}$ using a text-embedding model $f_{embed}(\cdot)$, where $d$ is the embedding dimension.
For each original prompt $x_{k}\in\mathcal{X}$ (with $k\in\{1,\dots,K\}$), we identify a set of political phrases $\mathcal{E}_{k}=\mathrm{Extract}(x_{k})$, where $\mathrm{Extract}(\cdot)$ returns political keywords (political figures or objects). We embed each extracted keyword $e\in\mathcal{E}_{k}$ as $\mathbf{u}_{k,e}=f_{embed}(e)\in\mathbb{R}^{d}$. For a given input $x_{k}$, we compute a country-level geopolitical sensitivity score by taking the maximum cosine similarity between any keyword embedding and any paragraph embedding for that country:
(4) $\hat{s}^{cc}_{k,i}=\max_{e\in\mathcal{E}_{k}}\;\max_{p\in\mathcal{P}_{i}}\frac{\mathbf{u}_{k,e}^{\top}\mathbf{v}_{i,p}}{\lVert\mathbf{u}_{k,e}\rVert\,\lVert\mathbf{v}_{i,p}\rVert}.$
This max-over-(keywords, paragraphs) operation implements a worst-case assumption: if any paragraph in the country’s Wikipedia description strongly aligns with any political keyword from the input, the input is treated as having a strong association with that country’s historical or sociopolitical context.
@@
This aggregation again follows worst-case reasoning. If any country within the same primary-language group exhibits a strong association with the input, the language-group score should reflect that highest potential sensitivity. The resulting output for input $k$ is the country common knowledge–based geopolitical sensitivity score:
(6) $\mathbf{S}_{cc}(k,\ell)=\left(S^{cc}_{k,\ell}\right)_{\ell\in\mathcal{L}}\in\mathbb{R}^{|\mathcal{L}|},$
which summarizes the maximum country-associated common-knowledge sensitivity of the input across all language groups.
3.4.3. bias-based
Unlike the preceding metrics, which evaluate political sensitivity using external knowledge sources, the bias-based metric directly assesses bias propagated into the IPDM descriptions themselves. The objective of this metric is to prevent bias associated with the original controversial term from being transferred into the IPDM description, while preserving the intended semantics of the target concept. Let $k$ denote the politically sensitive keyword and let $\tau(k)$ be its representative-language realization (defined in Section 3.2). For each language $\ell\in\mathcal{L}$, $PC^{2}$ produces an IPDM description paragraph $a_{e,\ell}$. We embed the representative term and the IPDM description as $\mathbf{u}_{k}=f_{embed}(\tau(k))\in\mathbb{R}^{d}$ and $\mathbf{v}_{k,\ell}=f_{embed}(a_{k,\ell})\in\mathbb{R}^{d}$, respectively.
@@
(8) $\mathbf{S}_{bb}(k,\ell)=(S^{bb}_{k,\ell})_{\ell\in\mathcal{L}}\in\mathbb{R}^{|\mathcal{L}|}.$
3.4.4. politics-based
@@
This similarity score serves as an indicator of political semantic proximity. Lower values indicate that the translated IPDM description is less semantically aligned with political concepts, which is desirable for our purpose. By selecting languages with lower politics-based scores, $PC^{2}$ prefers translations that convey the intended meaning while comparatively reducing political semantic associations. The resulting output is the language-wise politics-based score vector
(10) $\mathbf{S}_{pb}(k,\ell)=(S^{pb}_{e,\ell})_{\ell\in\mathcal{L}}\in\mathbb{R}^{|\mathcal{L}|}.$
3.5. Combined Score Aggregation
@@
For each target and candidate language that passes the back-translation filter, $PC^{2}$ computes a unified score by aggregating the individual metric scores described in Section 3.3. The combined score is computed as a weighted sum of these metric scores and is defined as follows:
$S_{\text{combined}}(k,\ell)={}$ $w_{kc}\,\mathbf{S}_{kc}(k,\ell)+w_{cc}\,\mathbf{S}_{cc}(k,\ell)$
${}+w_{bb}\,\mathbf{S}_{bb}(k,\ell)+w_{pb}\,\mathbf{S}_{pb}(k,\ell).$
Here, $\mathbf{S}_{kc}(k,\ell)$, $\mathbf{S}_{cc}(k,\ell)$, $\mathbf{S}_{bb}(k,\ell)$, and $\mathbf{S}_{pb}(k,\ell)$ denote the keyword common knowledge–based, country common knowledge–based, bias-based, and politics-based scores for language $\ell$, respectively, and $w_{kc},w_{cc},w_{bb},w_{pb}\geq 0$ control their contributions. $k$ denotes politically sensitive keyword.
@@ 3.6. Adversarial Prompt Construction
Using the combined scores computed in Section 3.5, $PC^{2}$ constructs the final adversarial prompt through a structured, model-aware selection and assembly process. For each target, candidate languages are sorted according to their combined scores. Rather than always selecting the top-ranked candidate, $PC^{2}$ employs an index-based strategy over the sorted list to balance semantic correctness and generation success, depending on the sensitivity of the target model. To determine appropriate selection indices, we conduct a bin-wise evaluation at the 0th, 25th, 50th, and 75th percentiles of the sorted candidate list using 60 samples. This analysis allows us to empirically characterize the trade-off between correctness and success rate and to select indices that are well-suited to different model behaviors.
After selecting the final language for each target, $PC^{2}$ assigns an alphabetical index to each IPDM description (e.g., A, B, C). The IPDM descriptions are then presented as an indexed list, such as A: <IPDM description 1> and B: <IPDM description 2>. The base prompt contains placeholders corresponding to the detected controversial terms, and each placeholder is replaced with its assigned alphabetical index, ensuring a consistent reference between the base prompt and the IPDM descriptions. Finally, the adversarial prompt is constructed by concatenating the indexed IPDM description list with the instantiated base prompt. This indirection enables the target model to infer the intended concepts through association rather than explicit mention, completing the adversarial prompt construction.
Because the final adversarial prompt may combine IPDM descriptions expressed in different languages, textual phrases appearing in the generated images may initially be rendered in multiple languages. In our experiments, we observe that such phrases can be converted to English when explicitly requested (e.g., "Convert the phrase in the placard to English"), and we inspect the resulting images accordingly. By intuition, similar phrases could be translated into other languages as well; however, we do not further investigate this aspect, as English is widely used as a common language in global contexts and provides a practical reference for consistent evaluation.

### v2

@@ 3. $PC^{2}$ Design
3.1. Politically Sensitive Keyword Detection
@@
Politically sensitive term detection is performed through a multi-stage pipeline with two independent processing paths. First, named entity recognition (NER) is applied to the input prompt to identify political figures (e.g., Donald Trump in Figure 5). For each identified political figure, a dedicated inference step based on a language model is used to determine the most relevant country. Separately, $PC^{2}$ extracts noun phrases that may be associated with a specific country or with politically or socially sensitive topics (e.g., Al-Qaeda flag in Figure 5). The extracted noun phrases are then analyzed using a language model (GPT-4o in our implementation) to determine their association with a specific country and their political or social sensitivity. This design enables the identification of politically sensitive terms beyond surface-level keyword matching. The inferred country information from both processing paths is retained for subsequent metric computation and is not used to modify the prompt at this stage.
3.2. Identity Preserving Descriptive Mapping
After identifying sensitive keywords, $PC^{2}$ performs Identity Preserving Descriptive Mapping (IPDM) for each keyword. As shown in Figure 5, this process replaces explicit sensitive keywords with nuanced descriptions that encapsulate the historical, political, and visual essence of the target political entities. For example, the keyword Donald Trump is replaced with several sentences describing his background (the blue box). For each sensitive keyword, $PC^{2}$ employs a dedicated language model (gpt-4o), generating 1–2 sentence quiz-style descriptions that implicitly convey the core concept without triggering prohibited terms (see Appendix B). The implementation of IPDM serves two critical strategic objectives: First, it effectively bypasses keyword-based pre-filters. Since these filters rely on direct matching of explicit sensitive terms (e.g., Al-Qaeda), IPDM’s descriptive approach allows the underlying intent to pass through undetected by stripping away the lexical triggers. Second, it provides the image generation model with sufficient descriptive cues to reconstruct the target entity accurately. By supplying rich details instead of isolated labels, IPDM ensures that the model can maintain the identity preservation required for high-quality PSCs.
3.3. Geopolitical Translation
@@
There are two types of targets that require the selection of a final language to construct the adversarial prompt: the IPDM descriptions and the base prompt. The IPDM descriptions are generated in the previous step (IPDM), while the base prompt corresponds to the original prompt in which controversial terms are replaced with placeholders (i.e., <PCC_PLACE_HOLDER_x>). These placeholders are later substituted with alphabetical indices (e.g., A, B, etc.), each mapped to its corresponding IPDM description. Both targets are initially translated into 72 languages (Appendix A).
@@
The current implementation of $PC^{2}$ uses GPT-4o as the translation model. This choice enables the collection of language-specific statistics of GPT models for evaluation (Section 5.1); however, the methodology is not model-specific, and any high-performing translation model could be used in its place. Nevertheless, even high-performing models may produce hallucinations or translation errors, particularly for low-resource languages. To mitigate this issue, we apply a back-translation procedure. Each translated output is translated back into English (e.g., English → Thai → English) and compared with the original content using cosine similarity. $PC^{2}$ filters out translations whose back-translated similarity score falls below an acceptable threshold (cosine similarity <0.9<0.9), thereby removing unreliable translations from further consideration.
3.4. Geopolitical Sensitivity Metrics
@@
To select an appropriate language for each translation target within a prompt, including both IPDM descriptions and the base prompt, we devise four metrics. These metrics assess the potential level of political sensitivity in the translated output for each language and assign a corresponding score. While the metrics are primarily defined over political keywords and thus applied to IPDM descriptions, we determine the language for the base prompt by computing the cumulative sum of the language scores obtained from the associated IPDM descriptions. This design choice is justified because the base prompt is intentionally neutral, containing only placeholders combined with action verbs (e.g., <PCC_PLACE_HOLDER_1> holding <PCC_PLACE_HOLDER_2>).
@@ 3.4.1. Keyword Common Knowledge-based Metric
The keyword common knowledge–based metric measures how controversially a political input may be perceived, using publicly available historical associations between the input’s keywords and countries, then aggregating those signals into geopolitical sensitivity score.
To evaluate an IPDM description using this metric, we first obtain its associated political keyword $k$ from the input prompt. By using this political keyword, we retrieve relevant Wikipedia paragraphs using $\mathrm{WikiSearch}(\cdot)$ and merge all retrieved content into a single paragraph set $\mathcal{P}_{k}=\bigcup\mathrm{WikiSearch}(k)$. The retrieved Wikipedia text is stored as paragraphs, which serve as the atomic retrieval units in a knowledge database. This database can be dynamically expanded as new phrases are detected in subsequent inputs. Each paragraph $p\in\mathcal{P}_{k}$ is encoded into a dense vector representation $\mathbf{v}_{k,p}=f_{embed}(p)\in\mathbb{R}^{d}$ using a text-embedding model $f_{embed}(\cdot)$, where $d$ is the embedding dimension.
To measure political sensitivity with respect to geopolitical actors, we construct a country-specific prompt by concatenating a fixed prefix with the country name. For example, country-specific prompt $q_{i}=\mathrm{Conf}\|\mathrm{name}_{i}$, where $\mathrm{Conf}$ is the string “Conflict with” and $\mathrm{name}_{i}$ denotes the name of country $i\in\{1,\dots,N\}$. We embed each country-specific prompt as $\mathbf{u}_{i}=f_{embed}(q_{i})\in\mathbb{R}^{d}$. For a given political keyword $k$, we compute a country-level geopolitical sensitivity score by taking the maximum cosine similarity between the country prompt embedding and any evidence paragraph embedding retrieved for that input:
@@
(3) $\mathbf{S}_{kc}(k)=\left(S^{kc}_{k,\ell}\right)_{\ell\in\mathcal{L}}\in\mathbb{R}^{|\mathcal{L}|},$
@@ 3.4.2. Country Common Knowledge-based Metric
We first enumerate a set of countries $\mathcal{C}=\{1,\dots,N\}$ and retrieve relevant Wikipedia paragraphs for each country name $\mathrm{name}_{i}$ using $\mathrm{WikiSearch}(\cdot)$. We merge the retrieved content into a country-specific paragraph set $\mathcal{P}_{i}=\mathrm{WikiSearch}(\mathrm{name}_{i})$. The collected Wikipedia text is segmented into paragraphs, which serve as the atomic retrieval units in a knowledge database. This database can be dynamically expanded as new countries are added. Each paragraph $p\in\mathcal{P}_{i}$ is encoded into a dense vector representation $\mathbf{v}_{i,p}=f_{embed}(p)\in\mathbb{R}^{d}$ using a text-embedding model $f_{embed}(\cdot)$, where $d$ is the embedding dimension.
Then, we embed the political keyword $k$ of the target IPDM description as $\mathbf{u}_{k}=f_{embed}(k)\in\mathbb{R}^{d}$. For a political keyword $k$ and country $i$, we compute a country-level geopolitical sensitivity score by taking the maximum cosine similarity between the keyword embedding and any paragraph embedding for that country:
(4) $\hat{s}^{cc}_{k,i}=\max_{p\in\mathcal{P}_{i}}\frac{\mathbf{u}_{k}^{\top}\mathbf{v}_{i,p}}{\lVert\mathbf{u}_{k}\rVert\,\lVert\mathbf{v}_{i,p}\rVert}.$
This max-over-paragraphs operation implements a worst-case assumption: if any paragraph in the country’s Wikipedia description aligns with the input, the input is treated as having a strong association with that country’s historical or sociopolitical context.
@@
This aggregation again follows worst-case reasoning. If any country within the same primary-language group exhibits a strong association with the input, the language-group score should reflect that highest sensitivity. The resulting output for the keyword $k$ is the country common knowledge–based geopolitical sensitivity score:
(6) $\mathbf{S}_{cc}(k)=\left(S^{cc}_{k,\ell}\right)_{\ell\in\mathcal{L}}\in\mathbb{R}^{|\mathcal{L}|},$
which summarizes the maximum country-associated common knowledge sensitivity of the input across all language groups.
3.4.3. Bias-based Metric.
Unlike the preceding metrics, which evaluate political sensitivity using external knowledge sources, the bias-based metric directly assesses bias propagated into the IPDM descriptions themselves. The objective of this metric is to prevent bias associated with the original political term from being transferred into the IPDM description, while preserving the intended semantics of the target concept. Let $k$ denote the politically sensitive keyword and let $\tau(k)$ be its representative-language realization (defined in Section 3.2). For each language $\ell\in\mathcal{L}$, $PC^{2}$ produces an IPDM description paragraph $a_{e,\ell}$. We embed the associated political term and the IPDM description as $\mathbf{u}_{k}=f_{embed}(\tau(k))\in\mathbb{R}^{d}$ and $\mathbf{v}_{k,\ell}=f_{embed}(a_{k,\ell})\in\mathbb{R}^{d}$, respectively.
@@
(8) $\mathbf{S}_{bb}(k)=(S^{bb}_{k,\ell})_{\ell\in\mathcal{L}}\in\mathbb{R}^{|\mathcal{L}|}.$
3.4.4. Politics-based Metric.
@@
This similarity score serves as an indicator of political semantic proximity. Lower values indicate that the translated IPDM description is less semantically aligned with political concepts, which is desirable for our purpose. By selecting languages with lower politics-based scores, $PC^{2}$ prefers translations that convey the intended meaning while comparatively reducing political semantic associations. The resulting output is the language-wise politics-based score vector:
(10) $\mathbf{S}_{pb}(k)=(S^{pb}_{e,\ell})_{\ell\in\mathcal{L}}\in\mathbb{R}^{|\mathcal{L}|}.$
3.4.5. Combined Score Aggregation
@@
For each target and candidate language that passes the back-translation filter, $PC^{2}$ computes a geopolitical sensitive score by aggregating the individual metric scores described in Section 3.3. The combined score is computed as a weighted sum of these metric scores and is defined as follows:
$S_{\text{combined}}(k)={}$ $w_{kc}\,\mathbf{S}_{kc}(k)+w_{cc}\,\mathbf{S}_{cc}(k)$
${}+w_{bb}\,\mathbf{S}_{bb}(k)+w_{pb}\,\mathbf{S}_{pb}(k).$
Here, $\mathbf{S}_{kc}(k)$, $\mathbf{S}_{cc}(k)$, $\mathbf{S}_{bb}(k)$, and $\mathbf{S}_{pb}(k)$ denote the keyword common knowledge–based, country common knowledge–based, bias-based, and politics-based scores of political keyword $k$ accross all language groups, respectively, and $w_{kc},w_{cc},w_{bb},w_{pb}\geq 0$ control their contributions.
@@ 3.5. Adversarial Prompt Construction
Using the combined scores computed in Section 3.4.5, $PC^{2}$ constructs the adversarial prompt through model-aware selection and assembly process. For each target, candidate languages are sorted according to their combined scores. Rather than selecting the top-ranked candidate, $PC^{2}$ employs an index-based strategy over the sorted list to balance semantic correctness and generation success, depending on the sensitivity of the target model. To determine appropriate selection indices, we conduct a bin-wise evaluation at the 0th, 25th, 50th, and 75th percentiles of the sorted candidate list using 60 samples. This analysis allows us to empirically characterize the trade-off between correctness and success rate and to select indices that are well-suited to different model behaviors.
After selecting the language for each target, $PC^{2}$ assigns an alphabetical index to each IPDM description (e.g., A, B, C). The IPDM descriptions are then presented as an indexed list, such as A: <IPDM description 1> and B: <IPDM description 2>. The base prompt contains placeholders corresponding to the detected controversial terms, and each placeholder is replaced with its assigned alphabetical index, ensuring a reference between the base prompt and the IPDM descriptions. Finally, the adversarial prompt is constructed by concatenating the indexed IPDM description list with the instantiated base prompt. This indirection enables the target model to infer the intended concepts through association rather than explicit mention, completing the adversarial prompt construction.
Because the final adversarial prompt may combine IPDM descriptions expressed in different languages, textual phrases appearing in the generated images may initially be rendered in multiple languages. In our experiments, we observe that such phrases can be converted to English when explicitly requested (e.g., "Convert the phrase in the placard to English"), and we inspect the images accordingly. Intuitively, similar phrases could be translated into other languages as well; however, we do not further investigate this, as English is widely used as a common language in global contexts and provides a practical reference for consistent evaluation.